Please help, last question was not answered right.

Using the product cost concept, determine the selling price for 30,000 units using the following data: product cost per unit, $15.00; total period costs, $90,000; and desired profit, $150,000. Round the markup percentage to one decimal place and markup to nearest dollar.

Total Product Costs = $Choose...23.0015.007.008.0010.00 x Choose...25,00020,00030,000 = $Choose...150,000120,000450,00090,000250,000

Markup Percentage = (Desired Profit + Total Period Costs) / Total Product Costs

= ($Choose...150,000120,000450,00090,000250,000 + $Choose...150,000120,000450,00090,000250,000 ) / $Choose...150,000120,000450,00090,000250,000

= Choose...53.3%42.0%20.0%33.3%

Mark up = $Choose...23.0015.007.008.0010.00 x Choose...53.3%42.0%20.0%33.3% = $Choose...23.0015.007.008.0010.00

Selling Price = $Choose...23.0015.007.008.0010.00 + $Choose...23.0015.007.008.0010.00 = $Choose...23.0015.007.008.0010.00   
Answer #1

Total Product Costs = 30,000 x 15

= $450,000

Markup Percentage = (Desired Profit + Total Period Costs) / Total Product Costs

= (150,000 + 90,000)/450,000

= 240,000/450,000

= 53.3%

Mark up = 450,000 x 53.3%

= $239,850

Mark up for 1 unit = 239,850/30,000

= $8

Selling price per unit = Product cost + Mark up

= 15 + 8

= $23

Kindly comment if you need further assistance. Thanks